# A measure-valued HJB perspective on Bayesian optimal adaptive control

### v1

@@ 2.1. Controls and objective
$\mathcal{U}\subseteq\left\{v=(v_{1},v_{2},\dots):\textrm{$v_{i}\in\mathbb{R}$, $i\in\mathbb{N}$, and $\sum_{i=1}^{\infty}(R^{i}v_{i})^{2}\leq K$}\right\};$
we here consider the subspace topology inherited from the product topology and note that $\mathcal{U}$ thus defined is a Polish space. We also define
(2) $h(v,x)=\sum_{i=1}^{\infty}v_{i}x^{i},\quad v\in\mathcal{U},\;x\in(-R,R);$
we note that this function is well-defined and jointly measurable, that the definition fixes $h(v,0)\equiv 0$, and that $x\mapsto h(v,x)$ is continuous uniformly in $v$ on $\operatorname{supp}(\mu_{0})$ (c.f. Lemma 4.1 below).
@@
[…]
@@
[…]
for further discussion on how this type of closed-loop controls relate to alternative definitions, we refer to [16, Remark 2.4].
@@ 2.2. MVMs and SDEs
For $u\in\mathcal{A}$, it holds that $\pi^{u}(f)$ is a martingale for any $f\in C_{b}$; $\pi^{u}$ is therefore a so-called measure-valued martingale (MVM) (see e.g. [17, Definition 2.1] or [5]; c.f. also e.g. [19] for an earlier account). More pertinently, taking the signal to be constant in the classical (controlled) Kushner-Stratonovich equation (see e.g. [43, Proposition 7.2.8]), we have that $\pi^{u}$ is the solution to 222 A-priori, setting $\mathcal{L}\equiv 0$ in [43, Proposition 7.2.8], we obtain (8) for $\varphi\in C^{2}_{b}$ with all of its derivatives bounded. However, for any $\varphi\in C_{b}$, there exists a sequence $\{\varphi_{n}\}$ belonging to this class and converging pointwise to $\varphi$. Since, by Lemma 4.1 below, $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$, $\mathbb{E}\int_{0}^{t}\left(\int_{\mathbb{R}}(\varphi-\varphi_{n})(x)\left(h(u_{s},x)-\pi^{u}_{s}\left(h(u_{s},\cdot)\right)\right)\pi^{u}_{s}(\mathrm{d}x)\right)^{2}\mathrm{d}s\xrightarrow[n\to\infty]{}0,$ which implies that (8) holds for each $\varphi\in C_{b}$.
@@
The links to MVMs and the above equations will become crucial for our upcoming analysis. Indeed, the benefit of linking our problem to equation (8) is that it enables relating the problem to a more standard (albeit still infinite-dimensional) stochastic control problem featuring a controlled state-process being driven by a Brownian motion; the idea being that if one manages to construct a solution to that auxiliary problem, one should be able to construct a solution to the original problem too. To see this, suppose for illustrative purposes that we’ve been given some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$ supporting a Brownian motion $B$, and have found a control in feedback form provided by a function $\hat{u}:\mathcal{P}\to\mathcal{U}$ for which the SDE
(9) $\mathrm{d}\tilde{\pi}_{t}(\varphi)$ $=\left(\tilde{\pi}_{t}\left(\varphi(\cdot)h\left(\hat{u}(\tilde{\pi}_{t}),\cdot\right)\right)-\tilde{\pi}_{t}\left(\varphi\right)\tilde{\pi}_{t}\left(h\left(\hat{u}(\tilde{\pi}_{t}),\cdot\right)\right)\right)\mathrm{d}B_{t},\quad\tilde{\pi}_{0}=\mu_{0},$
@@
admits a pathwise unique solution $\tilde{\pi}$. Then one would hope to be able to construct333 Indeed, returning to equations (3) and (6) with $u_{t}$ replaced by $\hat{u}(\pi^{u}_{t})$ in all instances, given that $\hat{u}$ is nice enough, these SDEs admit a pathwise unique solution, say $(Y,\pi)$, which is adapted to the filtration generated by $Y$. Defining $u^{*}\in\mathcal{A}^{pre}$ by $u^{*}_{t}=\hat{u}(\pi_{t})$, (3) gives that $Y=Y^{u^{*}}$ and it follows that the pair $(Y,\pi)$ is adapted to $\mathcal{Y}^{u^{*}}$; hence $u^{*}\in\mathcal{A}$. Since the solution to (6) is pathwise unique, we also have that $\pi=\pi^{u^{*}}$ and therefore $u^{*}_{t}=\hat{u}(\pi^{u^{*}}_{t})$. In consequence, $(Y^{u^{*}},\pi^{u^{*}})$ must solve the system of SDEs (7) and (8) with $u_{t}$ replaced by $\hat{u}(\pi^{u^{*}}_{t})$ and $I^{u}$ by $I^{u^{*}}$ which is a $(\mathcal{Y}^{u^{*}},\mathbb{P})$-BM. Since pathwise uniqueness implies uniqueness in law for the SDE (9), or equivalently for (8), evaluating the objective at $(\pi^{u^{*}},u^{*})$ and $(\tilde{\pi},\hat{u}(\tilde{\pi}_{\cdot}))$ must yield the same value, which verifies the claim. a control $u^{*}\in\mathcal{A}$ for which the objective attains the same value as the objective evaluated at the tuple $(\tilde{\pi},\hat{u}(\tilde{\pi}_{\cdot}))$. Provided (i) that $\hat{u}$ is optimal for the problem of optimising the objective over solutions to (9) for some suitable class of controls, and (ii) that the value of this problem coincides with the value of our original problem of interest, it would follow that $u^{*}$ were optimal for that original problem. In the upcoming analysis, the role of $\hat{u}$ will be played by piecewise constant controls in feedback form on a finite time grid, and we will establish that the value of the respective problem formulations coincide by linking them to the unique continuous viscosity solution of an associated HJB equation.
2.3. Related equations in the literature
Within the literature on stochastic control under partial information the above approach of linking the problem of original interest to a more standard control problem (involving measure-valued processes) is often referred to as the separation principle (although there is some ambiguity in how this terminology is used). We here recall the various related equations which have emerged within this context and review the related literature.
To this end, recall that one is here typically interested in a problem which is similar to ours but where the hidden signal follows itself an Itô dynamics; specifically, one is allowed to control that signal process, say $X_{t}$, $t\geq 0$, while the observation is given by (3) for some fixed (not controlled [^4] Notably, [43] and [2] also allow for controlled observations.) function $h(x)$, $x\in\mathbb{R}$, and $X$ is replaced by $X_{t}$. The aim is to optimise an objective of the form (1) — again with $X$ replaced by the controlled process $X_{t}$. The nomenclature ‘partially observable’ refers to the fact that the controls must be adapted to the filtration generated by the observation process; controls given in feedback form in terms of the signal are for example not admissible. As discussed in the introduction, one often considers wide-sense controls that are adapted to a somewhat larger filtration; it is also common to consider various weak formulations where the explicit control-dependence of the filtration is circumvented.
Kushner-Stratonovich equations
In its original form, the separation principle amounts to rewrite the objective onto the form (5), where now
$\pi^{u}_{t}=\mathrm{Law}(X_{t}|\mathcal{Y}^{u}_{t});$
this filter then solves the so-called Kushner-Stratonovich equation which is similar to (8) but with some additional terms appearing due to the dynamics of $X$ itself — the control then appears through those additional terms while typically $h(v,x)\equiv h(x)$. At the cost of the involved processes taking values in the space of probability measures, the resulting problem of optimising (5) over solutions to the KS equation is conceptually simpler since it is a fully observable problem; for example, controls given in feedback form in terms of the filter are admissible. Provided an optimiser can be found, one would aim at closing the loop by constructing a solution to the original problem along the lines outlined above.
While offering a conceptually simple and elegant formulation, this problem turns out to be notoriously difficult to solve in practice, and explicit results are few to date. Particular attention has however been paid to the case when $X_{0}$ is Gaussian and the various coefficients appearing in the dynamics of the signal and the observation process are linear. The resulting filter, referred to as the Kalman-Bucy filter, is then Gaussian too with its variance process solving a deterministic Ricatti equation and its mean process an SDE. The separated problem thus becomes a finite-dimensional problem which can be addressed using standard methods. For various results in this direction we refer to [26], [9, Chapter 7] and [43, Section 7.5]; see in particular [43, Proposition 7.3.9] for a result analogous to the above description of how to construct an optimiser to the original problem based on an optimiser to the separated one.
Zakai equations
A more common approach has been to employ the separation principle in conjunction with the so-called Zakai equation. The approach is based on introducing a measure $\mathbb{Q}^{u}_{t}\sim\mathbb{P}$ under which $Y^{u}$ is a Brownian motion on $[0,t]$ (independent of $X_{0}$ and the Brownian motion driving the signal). By Girsanov, this is achieved by defining d/$\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}^{u}_{t}=\Lambda^{u}_{t}$ with $\mathrm{d}\Lambda^{u}_{t}=h(X_{t},u_{t})\Lambda^{u}_{t}\mathrm{d}Y^{u}_{t}$. Defining the measure-valued process
(10) $\sigma^{u}_{t}(\varphi):=\mathbb{E}^{\mathbb{Q}^{u}_{t}}[\varphi(X_{t})\Lambda^{u}_{t}|\mathcal{Y}^{u}_{t}],\quad\varphi\in C_{b},$
making a change of measure, and, in turn, applying the tower property, one may then rewrite the objective as follows
(11) $\mathbb{E}^{\mathbb{Q}^{u}_{T}}\left[\int_{0}^{\infty}e^{-\beta t}\sigma^{u}_{t}\left(\tilde{k}(\cdot,u_{t})\right)\mathrm{d}t\right];$
meanwhile, making use of the properties of $\mathbb{Q}^{u}_{t}$, one obtains that the process $\sigma^{u}$ solves the so-called Zakai equation. For our set-up with a constant signal and controlled observations, this equation would take the form
(12) $\mathrm{d}\sigma^{u}_{t}(\varphi)=\sigma^{u}_{t}\left(h(u_{t},\cdot)\varphi\right)\mathrm{d}Y^{u}_{t}.$
For the case with a signal following a controlled Itô dynamics there are however additional terms present; the control notably enters through those terms and typically $h(v,x)\equiv h(x)$. The process $\sigma^{u}$ is referred to as the unnormalised process: by Bayes’ rule, $\pi^{u}_{t}(\varphi)=\sigma^{u}_{t}(\varphi)/\sigma^{u}_{t}(1)$, $\varphi\in C_{b}$.
The main point is that one may again view the problem of optimising (11) over controlled solutions to the Zakai equation (c.f. (12)) as a fully observed stochastic control problem (with a measure-valued controlled process). To study this separated problem has been the predominant approach in the literature. Most frequently, it is assumed that the unnormalised conditional distribution admits a density: the Zakai equation then reduces to the so-called Duncan-Mortensen-Zakai equation for the density process itself. Within such a set-up, there are numerous contributions to date: In [9, Chapter 8] the problem is dealt with by use of the stochastic maximum principle as well as via the DPP combined with semi group arguments. In [32] (c.f. [30, 31]) the associated HJB-equation is studied and it is proven that the value function is indeed a viscosity solution (albeit in a rather weak sense which does not enable getting comparison). In [23], the question of existence of optimal controls is discussed (within the class of randomised controls). In [38], it is assumed that the density process belongs to $L^{2}$; it is then established that the value function is indeed the unique viscosity solution to the corresponding HJB equation (see also [37, 39]). In [28], the weaker requirement that the density process belongs to some weighted $L^{2}$-space is imposed (still a Hilbert space); also here, it is then proven that the value function is the unique viscosity solution to the HJB equation. Finally, motivated by similar questions, [40] established an Itô formula for solutions to the KS and Zakai equations without imposing any assumptions on the existence of densities; in [41], viscosity solutions were discussed for some related equations.
Yet another version of the separation principle
The articles [2] and [3] considered yet a different implementation of the separation principle. By first making a change of measure and using again the tower property, they noticed that the objective can also be rewritten as
(13) $\mathbb{E}^{\mathbb{Q}^{u}_{T}}\left[\int_{0}^{\infty}\int e^{-\beta t}\tilde{k}(x,u_{t})z\;\mathrm{d}\pi^{u}_{t}(x,z)\mathrm{d}t\right],$
where $\pi^{u}_{t}$ is now defined as the joint distribution of $(X_{t},\Lambda^{u}_{t})$ under $\mathbb{Q}^{u}_{t}$ given $\mathcal{Y}^{u}_{t}$. Recalling the dynamics for $\Lambda_{t}$ and rewriting the dynamics for $X$ in terms of the observation process, one might then write down the equation for $\pi^{u}$.
Indeed, effectively one is then back in a set-up similar to the one considered when deriving the KS equation with the difference that the only thing that is being observed is some correlated noise (note that in this case the unnormalised and normalised conditional probabilities coincide); in effect, one has thus obtained a simpler dynamics at the cost of introducing an additional factor. For our case with a constant signal and controlled observations, the equation would take the form
$\mathrm{d}\pi^{u}_{t}(\varphi)=\frac{1}{2}\int\left(h(u_{t},x)z\right)^{2}\dfrac{\partial^{2}\varphi}{\partial z^{2}}(x,z)\mathrm{d}\pi^{u}_{t}(x,z)\mathrm{d}t+\int h(u_{t},x)z\dfrac{\partial\varphi}{\partial z}(x,z)\mathrm{d}\pi^{u}_{t}(x,z)\mathrm{d}Y^{u}_{t};$
for the general case there are however additional terms present. The main point is that one may now consider the separated problem where the objective is given by (13) and where the controlled probability-measure valued processes $\pi^{u}$ follows a certain dynamics.
In [3], this separated problem is studied in a Markovian set-up; the article notably imposes no assumptions on the existence of a density. They consider the corresponding problem using randomised controls and first establish that the randomised problem and the original problem coincide. For the randomised problem they then consider the HJB equation and prove that the value function is the unique viscosity solution. To do so, they make use of the Lions derivative; that is, they work in a lifted space and rely on the Itô formula in terms of the Lions derivative as formulated in e.g. [13]. In [2], the authors consider the analogous (separated) problem with non-Markovian dynamics and address the problem using BSDE methods.
Notably, [2] also allows for controlled observations. We leave for future research to investigate whether their approach may provide further insights compared to the results obtained herein.
A finite-dimensional reduction
@@ 2.3. Related equations in the literature
In [16], the authors consider a set-up more closely related to ours where the signal is constant and the effect of the control enters through the observation. Specifically, their observation process follows a dynamics similar to (3) but with $h(u_{t},X)$ replaced by $b(t,Y^{u}_{t},u_{t})X$ for some function $b$; they also let the signal be multi-dimensional and allow for a controlled volatility coefficient but we omit those features here. For this set-up, $\sigma^{u}$ is given by (10) with $X_{t}\equiv X$ and d/$\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}^{u}_{t}=\Lambda^{u}_{t}=\mathcal{E}(X\Upsilon^{u})_{t}$, where
@@
Hence, defining, $F[\varphi](v,\gamma)=\int\varphi(x)e^{xv-\frac{1}{2}x^{2}\gamma}\mu_{0}(\mathrm{d}x),$ it follows that
@@
In particular, applying Itô’s formula and making use of (14) and the fact that $\frac{1}{2}F[\varphi]_{vv}+F[\varphi]_{\gamma}=0$ and $F[\varphi]_{v}=F[\varphi\cdot\mathrm{id}]$, it is straightforward to verify that $\sigma^{u}$ thus defined satisfies
@@
which is what the Zakai equation reduces to in this set-up (c.f. (12)).
The filter is thus fully characterised by the current state of the processes $\Upsilon^{u}$ and $\langle\Upsilon^{u}\rangle$; in [16], this is exploited in order to address the problem of optimising, over $u\in\mathcal{A}$, the finite-horizon analogue of an objective of the form (1) — with the running cost $\tilde{k}(X,u_{t})$ replaced by $\tilde{k}(t,Y^{u}_{t},X,u_{t})$ — as a finite-dimensional problem. More pertinently, note that for $u\in\mathcal{A}$, the innovations process, which is a $(\mathcal{Y}^{u},\mathbb{P})$-BM, is given by
$\mathrm{d}I^{u}_{t}=\left(X-\pi^{u}_{t}(\mathrm{id})\right)b(t,Y^{u}_{t},u_{t})\mathrm{d}t+\mathrm{d}W_{t},\quad I^{u}_{0}=0;$
in consequence,
@@
Combining (14) with (16), they observe that the dynamics of $(Y^{u},\Upsilon^{u},\langle\Upsilon^{u}\rangle)$ can be described by a system of SDEs driven by a $\mathbb{P}$-BM. In parallel, they express the objective as an expected value under $\mathbb{P}$ of a function depending on precisely $(Y^{u},\Upsilon^{u},\langle\Upsilon^{u}\rangle)$; this can be achieved by rewriting the objective as an expected value of a functional depending on the observation and the filter (in analogy with the KS approach) and subsequently express the filter in terms of $(\Upsilon^{u},\langle\Upsilon^{u}\rangle)$. This leaves them with a separated problem in a more standard form, the study of which forms the basis for their approach.
As mentioned in [16, Remark 2.7], a similar approach may be employed when restricting to controls which are polynomials of some fixed degree; we crucially go beyond such an assumption in our work. It is worth noticing that when turning an optimiser of their separated problem into an optimiser for the original problem, they focus on piecewise constant controls; we will here follow a similar approach. In order to further relate their set-up to ours, making use of (15), (16) and Itô’s formula, we obtain that
$\mathrm{d}\pi^{u}_{t}(\varphi)=\left(\pi^{u}_{t}(\varphi\;\mathrm{id})-\pi^{u}_{t}(\varphi)\pi^{u}_{t}(\mathrm{id})\right)b(t,Y^{u}_{t},u_{t})\mathrm{d}I^{u}_{t}.$

### v2

@@ 2.1. Controls and objective
$\mathcal{U}\subseteq\left\{v=(v_{0},v_{1},\dots):\textrm{$v_{i}\in\mathbb{R}$, $i\in\mathbb{Z}_{+}$, and $\sum_{i=0}^{\infty}(R^{i}v_{i})^{2}\leq K$}\right\};$
we here consider the subspace topology inherited from the product topology and note that $\mathcal{U}$ thus defined is a compact Polish space. We let
(5) $h(v,x):=\sum_{i=0}^{\infty}v_{i}x^{i},\quad v\in\mathcal{U},\;x\in\operatorname{supp}(\mu_{0});$
we note that this function is well-defined and jointly measurable and that $x\mapsto h(v,x)$ is continuous and bounded on $\operatorname{supp}(\mu_{0})$ uniformly in $v$ (c.f. Lemma 4.1 below).
@@
$\mathcal{A}^{pre}:=\left\{\textrm{$u:\Omega\times[0,\infty)\to\mathcal{U}$: $u$ is $\mathbb{F}$-progressively measurable}\right\}.$
@@
$\mathcal{A}:=\left\{u\in\mathcal{A}^{pre}:\textrm{$u$ is $\mathcal{Y}^{u}$-progressively measurable}\right\};$
for further discussion on how this type of closed-loop controls relate to alternative definitions, we refer to [14, Remark 2.4].
@@
To further illustrate our setup, we give an example of a control problem which will fit into our setting, and for which our main results therefore are valid.
Example 2.2.
Consider the following scenario: player A seeks to determine the (unknown and stationary) location of player B, denoted by $X$, without revealing his own position. The location of player B is known to lie within I:=[1,1]I:=[-1,1], i.e. the $X$ is compactly supported on $I$. Player A is equipped with a sensor characterised by
(9) $\tilde{h}(\tilde{v},x)=e^{-\lambda(x-\tilde{v})^{2}},\quad 0<\lambda<\infty,\quad\tilde{v},x\in I,$
where $\tilde{v}$ denotes the location where player A places his sensor. The strength of the signal he receives is then determined by the distance between his sensor and the true location $x$, with the strength decaying exponentially as the squared distance increases.
Consequently, the information available to player A evolves according to the stochastic process
$\mathrm{d}Y^{\tilde{u}}_{t}=e^{-\lambda(X-\tilde{u}_{t})^{2}}\mathrm{d}t+\mathrm{d}W_{t},\quad Y^{\tilde{u}}_{0}=0,$
where $W$ is a standard Brownian motion independent of $X$, and where player A selects his position $\tilde{u}_{t}\in I$ at each time $t$ based on the observations collected up to that time. Player A begins with a prior distribution $\mu_{0}$ supported on $I$. The corresponding posterior distribution, or the filter, $\pi^{\tilde{u}}$ is defined as in (7). Player A then considers an objective similar to (8) with a running cost $\tilde{k}:\mathcal{P}\times I\to\mathbb{R}$ defined by
(10) $\tilde{k}(\mu,\tilde{v})=F\cdot\int_{[-1,1]}e^{-\theta(x-\tilde{v})^{2}}\mu(\mathrm{d}x)+G\cdot\mathbb{V}\text{ar}_{\mu}(\mathrm{id}),\quad F,\ G\geq 0,$
where $\mathbb{V}\text{ar}_{\mu}(\mathrm{id}):=\mu(\mathrm{id}^{2})-\mu(\mathrm{id})^{2}$, the coefficient $F$ represents the fear associated with exposure, and $G$ the cost to player A of uncertainty in their knowledge of the true location of player B; we assume player B has the same type of sensor as player A, possibly with a different sensitivity parameter $\theta>0$. Hence, the function $\tilde{k}$ encodes the trade-off between fear and reward: the first term penalises being too close to player B, while the second term rewards a reduction in uncertainty about player B’s location.
We anticipate a sensible expansion of the form $\tilde{h}(\tilde{v},x)=\sum_{i=0}^{\infty}C_{i}(\tilde{v})x^{i}$ allowing us to identify our set of controls as taking the form appropriate for the definitions above. In Appendix B, we rigorously verify such an expansion and show that under this reformulation, the problem satisfies all the conditions imposed above. It will then follow from our main result that we can characterise the value of this problem as the unique continuous viscosity solution of a measure-valued HJB equation, and construct an $\varepsilon$-optimal piecewise constant control in feedback form.
@@ 2.2. MVMs and SDEs
For $u\in\mathcal{A}$, it holds that $\pi^{u}(f)$ is a martingale in the observation filtration, for any $f\in C_{b}$, and $\pi^{u}$ is therefore a so-called measure-valued martingale (MVM); see [15, Definition 2.1] or [5], and also e.g. [17] for an earlier account. More pertinently, taking the signal to be constant in the classical (controlled) Kushner-Stratonovich equation (see e.g. [38, Proposition 7.2.8]), we have that $\pi^{u}$ is the solution to [^3] A-priori, setting $\mathcal{L}\equiv 0$ in [38, Proposition 7.2.8], we obtain (13) (or, equivalently, (11)) for $\varphi\in C^{2}_{b}$ with all of its derivatives bounded. However, for any $\varphi\in C_{b}$, there exists a sequence $\{\varphi_{n}\}$ belonging to this class and converging pointwise to $\varphi$. Since, by Lemma 4.1 below, $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$, $\mathbb{E}\int_{0}^{t}\left(\int_{\mathbb{R}}(\varphi-\varphi_{n})(x)\left(h(u_{s},x)-\pi^{u}_{s}\left(h(u_{s},\cdot)\right)\right)\pi^{u}_{s}(\mathrm{d}x)\right)^{2}\mathrm{d}s\xrightarrow[n\to\infty]{}0,$ which implies that (13) holds for each $\varphi\in C_{b}$.
@@
The links to MVMs and the above equations will become crucial for our upcoming analysis. Indeed, the benefit of linking our problem to equation (13) is that it enables relating the problem to a more standard (albeit infinite-dimensional) stochastic control problem featuring a controlled state-process being driven by a Brownian motion; the idea being that if one manages to construct a solution to this auxiliary (separated) problem, one should be able to construct a solution to the original problem too. To see this, suppose for illustrative purposes that we have found a control in feedback form provided by a function $\hat{u}:\mathcal{P}\to\mathcal{U}$ for which the SDE
(14) $\mathrm{d}\pi_{t}(\varphi)$ $=\left(\pi_{t}\left(\varphi(\cdot)h\left(\hat{u}(\pi_{t}),\cdot\right)\right)-\pi_{t}\left(\varphi\right)\pi_{t}\left(h\left(\hat{u}(\pi_{t}),\cdot\right)\right)\right)\mathrm{d}B_{t},\;\varphi\in C_{b},\;\;\pi_{0}=\mu_{0},$
@@
admits a pathwise unique solution. Provided (i) that $\hat{u}$ is optimal for the problem of optimising the objective over solutions to (14) for some suitable class of controls; (ii) that the value of this problem coincides with the value of our original problem of interest; and (iii) that by relying on uniqueness in law for (14), one may construct a control $u^{*}\in\mathcal{A}$ for which the objective attains the same value as when evaluated at the solution of this equation; then it would follow that this $u^{*}$ were optimal for the original problem. In the upcoming analysis, the role of $\hat{u}$ will be played by a piecewise constant control (in feedback form on a fine enough time grid) for which step (iii) simplifies considerably. In this way, we construct a piecewise constant and arbitrarily-close-to-optimal strategy for our main problem of interest. We will verify step (ii) by arguing that the value of the respective problems solve the same HJB equation; this will also give us further information about the value of the problem.
One question is whether the reduction to piecewise constant controls (which is an approach we take from [14]) is necessary, and whether one could work directly with a control derived in feedback form. However it appears to be more difficult to verify that an arbitrary feedback control will be adapted to the filtration which arises from implementing the control, and hence that the control is admissible. This has been shown in much simpler settings however, see e.g. [38, Proposition 7.3.9] for a result of this type in the much simplified Kalman-Bucy setting.
Remark 2.3 (A finite-dimensional reduction).
@@
In [14], the authors consider a setup closely related to ours: The dynamics of their observation process follows (6) with $h(u_{t},X)$ replaced by $b(t,Y^{u}_{t},u_{t})X$ for some function $b$ (they also let the signal be multi-dimensional and allow for a controlled volatility coefficient). For this setup, the unnormalised filter is given by $\sigma^{u}_{t}(\varphi)=\mathbb{E}^{\mathbb{Q}^{u}_{t}}[\varphi(X)\Lambda^{u}_{t}|\mathcal{Y}^{u}_{t}]$, $\varphi\in C_{b}$, where $\Lambda^{u}_{t}=\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}^{u}_{t}$ and $\mathbb{Q}^{u}_{t}$ is the measure rendering $Y^{u}$ a Brownian motion on $[0,t]$; that is, $\Lambda^{u}_{t}=\mathcal{E}(X\Upsilon^{u})_{t}$ with
@@
Defining $F[\varphi](v,\gamma)=\int\varphi(x)e^{xv-\frac{1}{2}x^{2}\gamma}\mu_{0}(\mathrm{d}x),$ it follows that
@@
Note that, by Itô’s formula,
@@
which is what the Zakai equation reduces to in this setup. Moreover, in this setting the innovations process is given by
$\mathrm{d}I^{u}_{t}=\left(X-\pi^{u}_{t}(\mathrm{id})\right)b(t,Y^{u}_{t},u_{t})\mathrm{d}t+\mathrm{d}W_{t},\quad I^{u}_{0}=0,$
and in consequence,
@@
Making use of (16), (17) and applying again Itô’s formula, it follows that
$\mathrm{d}\pi^{u}_{t}(\varphi)=\left(\pi^{u}_{t}(\varphi\;\mathrm{id})-\pi^{u}_{t}(\varphi)\pi^{u}_{t}(\mathrm{id})\right)b(t,Y^{u}_{t},u_{t})\mathrm{d}I^{u}_{t};$
@@
Crucially, the filter is here fully characterised by the current state of the processes $\Upsilon^{u}$ and $\langle\Upsilon^{u}\rangle$. In [14], this is carefully exploited: They express their objective — the finite-horizon analogue of (4) with $\hat{k}(X,u_{t})$ replaced by $\hat{k}(t,Y^{u}_{t},X,u_{t})$ — as an expected value (under $\mathbb{P}$) of a functional depending on $(Y^{u},\Upsilon^{u},\langle\Upsilon^{u}\rangle)$; combining (15) with (17), they then observe that the dynamics of $(Y^{u},\Upsilon^{u},\langle\Upsilon^{u}\rangle)$ can be described by a system of SDEs driven by a $\mathbb{P}$-Brownian motion. This leaves them with a finite-dimensional separated problem, the study of which forms the basis for their approach. As mentioned in [14, Remark 2.7], a similar approach may be employed when restricting to controls which are polynomials of some fixed degree; we crucially go beyond such an assumption.
Remark 2.4.
One might ask whether our results could be extended to include dependence on the observation process in the cost and observation functions. Including it in the cost $k$ means that the value function depends also on the factor $Y^{u}$; hence, one would need to verify that the viscosity and comparison theory we are citing from [16] carries over to this case. To include it in the observation function $h$, one would in addition need to establish existence of solutions to the pair of equations (12)–(13) when $h$ also depends on the $Y$-component of that solution (with $u$ being constant and $I^{u}$ some fixed Brownian motion; c.f. (20)) as well as study stability properties of that system of equations (c.f. Theorem 4.3). We leave it for future research to investigate those interesting questions.